) a particle of mass 40 micro gram and carrying a charge 5/1000000000C is moving directly towards a fixed positive charge of  magnitude 1/100000000C. when it is at a distance of 10cm from the fixed point charge, it has a velocity of 50cm/s. At what distance from the fixed charge will the particle comes to memontarily rest? is the acceleration constant during the motion?

Grade:12th pass

1 Answers

Eshan
5 years ago
When the charge comes to rest, it will lose all its kinetic energy to potential energy of the system.
Hence \dfrac{1}{2}mv^2+\dfrac{1}{4\pi\epsilon_0}\dfrac{q_1q_2}{r_1}=\dfrac{1}{4\pi\epsilon_0}\dfrac{q_1q_2}{r_2}
\implies \dfrac{1}{r_2}=\dfrac{1}{r_1}+\dfrac{2\pi\epsilon_0mv^2}{q_1q_2}
\implies r_2=8.25mm